# Effective Polaron Dynamics of an Impurity Particle Interacting with a Fermi Gas

### v1

@@ 3.2 Effective coherent state
$\leq$ $2\int_{0}^{t}\textnormal{d}s\int_{0}^{1}\textnormal{d}\tau\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}}\bigg{\{}\|\overline{(h_{y})_{\alpha}(k)}\mathcal{R}^{1-\tau}_{\alpha}(k)e^{\tau B}\psi\|+\|\mathcal{R}^{1-\tau}_{\alpha}(k)^{\ast}(h_{y})_{\alpha}(k)e^{\tau B}\psi\|\bigg{\}}$
$\leq$ $2\lambda\int_{0}^{t}\textnormal{d}s\int_{0}^{1}\textnormal{d}\tau\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}}|\hat{V}(k)n_{\alpha}(k)|\bigg{\{}\|\mathcal{R}^{1-\tau}_{\alpha}(k)e^{\tau B}\psi\|+\|\mathcal{R}^{1-\tau}_{\alpha}(k)^{\ast}e^{\tau B}\psi\|\bigg{\}}$
$\leq$ $C\int_{0}^{t}\textnormal{d}s\int_{0}^{1}\textnormal{d}\tau\ \lambda k_{\textnormal{F}}\|\hat{V}\|_{2}\bigg{\{}\bigg{(}\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}}\|\mathcal{R}^{1-\tau}_{\alpha}(k)e^{\tau B}\psi\|^{2}\bigg{)}^{1/2}+\bigg{(}\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}}\|\mathcal{R}^{1-\tau}_{\alpha}(k)^{\ast}e^{\tau B}\psi\|^{2}\bigg{)}^{1/2}\bigg{\}}$
@@
$\leq$ $C(f_{C}(\kappa k_{\textnormal{F}}t)-1)\lambda k_{\textnormal{F}}tMN^{-\frac{2}{3}+\delta}\|(\mathcal{N}+5)\psi\|.$ (5.22)
Again, using the choice of $\kappa=\lambda$ and $\psi=\phi\otimes\Omega$ we obtain the desired bound.
With the subsequent Lemma 5.4, we can conclude with a bound on $h_{0}$ of the form
@@
$\leq C\beta t\left\{\log\left[f_{1}(\kappa k_{\textnormal{F}}t)\right]+\log\left[f_{1}(\kappa k_{\textnormal{F}}t)\right]^{4}\right\}\left\{f_{C}(\kappa k_{\textnormal{F}}t)+1\right\}.$ (5.23)
@@
In total we expect, that all terms can be bounded by assumption on the initial condition on $\phi$. We will first focus on the term $\Delta_{y}W(\eta_{s})$. It holds
$W(\eta_{s})^{\ast}\partial_{y_{i}}W(\eta_{s})$ $=\left.e^{-\tau B_{s}}\partial_{y_{i}}e^{\tau B}\right|_{\tau=0}^{\tau=1}=\int_{0}^{1}\textnormal{d}\tau\ \partial_{\tau}\left(e^{-\tau B}\partial_{y_{i}}e^{\tau B}\right)$
$=\int_{0}^{1}\textnormal{d}\tau\left\{-Be^{-\tau B}\partial_{y_{i}}e^{\tau B}+e^{-\tau B}\partial_{y_{i}}\left(B_{s}e^{\tau B}\right)\right\}$
$=\int_{0}^{1}\textnormal{d}\tau\ e^{-\tau B}\left(\partial_{y_{i}}B_{s}\right)e^{\tau B}.$ (5.25)
With
@@
$\hphantom{=}+2i\int_{0}^{1}\textnormal{d}\tau\ W_{1-\tau}(\eta_{s})\text{Im}\langle ik_{i}\eta_{s},\mathcal{R}^{1-\tau}\rangle_{k,\alpha}W_{\tau}(\eta_{s}).$ (5.28)
@@
$\hphantom{=}+2i\int_{0}^{\tau}\textnormal{d}\sigma\ W_{1-\sigma}(\eta_{s})\text{Im}\langle ik_{i}\eta_{s},\mathcal{R}^{\tau-\sigma}\rangle_{k,\alpha}W_{\sigma}(\eta_{s})$ (5.29)
@@
$=$ $\left(c^{\ast}(k^{2}\eta_{s})-c(k^{2}\eta_{s})+i\text{Im}\langle\eta_{s},k^{2}\eta_{s}\rangle\right)W(\eta_{s})+2i\int_{0}^{1}\textnormal{d}\tau\ W_{1-\tau}(\eta_{s})\text{Im}\langle k^{2}\eta_{s},\mathcal{R}^{1-\tau}\rangle_{k,\alpha}W_{\tau}(\eta_{s})$
@@
$+2i\sum_{i=1}^{3}\int_{0}^{1}\textnormal{d}\tau\ \text{Im}\langle ik_{i}\eta_{s},\partial_{y_{i}}\left\{W_{1-\tau}(\eta_{s})\mathcal{R}^{1-\tau}_{\alpha}(k)W_{\tau}(\eta_{s})\right\}\rangle_{k,\alpha}$ (5.30)
$=$ $\left(c^{\ast}(k^{2}\eta_{s})-c(k^{2}\eta_{s})+i\text{Im}\langle\eta_{s},k^{2}\eta_{s}\rangle\right)W(\eta_{s})+2i\int_{0}^{1}\textnormal{d}\tau\ \text{Im}\langle k^{2}\eta_{s},\mathcal{R}^{1-\tau}\rangle_{k,\alpha}W_{\tau}(\eta_{s})$
@@
×$\times\left\{\left(c^{\ast}(ik_{i}\eta_{s})-c(ik_{i}\eta_{s})+i\text{Im}\langle\eta_{s},ik_{i}\eta_{s}\rangle\right)W(\eta_{s})+2i\int_{0}^{1}\textnormal{d}\tau\ \text{Im}\langle ik_{i}\eta_{s},\mathcal{R}^{1-\tau}\rangle_{k,\alpha}W_{\tau}(\eta_{s})\right\}$
$+2i\sum_{i=1}^{3}\int_{0}^{1}\textnormal{d}\tau\ \text{Im}\langle ik_{i}\eta_{s},\partial_{y_{i}}\left\{W_{1-\tau}(\eta_{s})\mathcal{R}^{1-\tau}_{\alpha}(k)W_{\tau}(\eta_{s})\right\}\rangle_{k,\alpha}$ (5.31)
@@
$I_{2}$ $\coloneqq 2i\int_{0}^{1}\textnormal{d}\tau\ W_{1-\tau}(\eta_{s})\text{Im}\langle k^{2}\eta_{s},\mathcal{R}^{1-\tau}\rangle_{k,\alpha}W_{\tau}(\eta_{s}),$ (5.33)
$I_{3}$ $\coloneqq 2\sum_{i=1}^{3}\left(c^{\ast}(ik_{i}\eta_{s})-c(ik_{i}\eta_{s})+i\text{Im}\langle\eta_{s},ik_{i}\eta_{s}\rangle\right)\int_{0}^{1}\textnormal{d}\tau\ iW_{1-\tau}(\eta_{s})\text{Im}\langle ik_{i}\eta_{s},\mathcal{R}^{1-\tau}\rangle_{k,\alpha}W_{\tau}(\eta_{s}),$ (5.34)
$I_{4}$ $\coloneqq 2i\sum_{i=1}^{3}\int_{0}^{1}\textnormal{d}\tau\ \text{Im}\langle ik_{i}\eta_{s},\partial_{y_{i}}\left\{W_{1-\tau}(\eta_{s})\mathcal{R}^{1-\tau}_{\alpha}(k)W_{\tau}(\eta_{s})\right\}\rangle_{k,\alpha}.$ (5.35)
@@
For $I_{1}$, we can treat all $c^{\ast}(\cdots)$ and $c(\cdots)$ terms with Lemma 5.1. Furthermore we use that $c_{\alpha}(k)\mathcal{N}=(\mathcal{N}+2)c_{\alpha}(k)$ to estimate
@@
For $I_{3}$, we first observe that for $n\in\mathbb{N}$
$\|\mathcal{N}^{n}\mathcal{R}^{\mu}_{\beta}(k)\psi\|$ $\leq\int_{0}^{\mu}\textnormal{d}\tau\ \|\mathcal{N}^{n}e^{-\tau B}\left(\sum_{l\in\Gamma^{\textnormal{nor}}}\eta_{\beta}(l)\mathcal{E}_{\beta}(l,k)\right)e^{\tau B}\psi\|$
$\leq\int_{0}^{\mu}\textnormal{d}\tau\ e^{C\|\eta_{s}\|\tau}\sum_{l\in\Gamma^{\textnormal{nor}}}|\eta_{\beta}(l)|\ \|\mathcal{E}_{\beta}(l,k)(\mathcal{N}+3)^{n}e^{\tau B}\psi\|$
$\leq C\int_{0}^{\mu}\textnormal{d}\tau\ e^{C\|\eta_{s}\|\tau}\sum_{l\in\Gamma^{\textnormal{nor}}}|\eta_{\beta}(l)|\ MN^{-\frac{2}{3}+\delta}\|\mathcal{N}(\mathcal{N}+3)^{n}e^{\tau B}\psi\|$
$\leq C\|\eta_{s}\|^{-1}(e^{C\|\eta_{s}\|\mu}-1)MN^{-\frac{2}{3}+\delta}\sum_{l\in\Gamma^{\textnormal{nor}}}|\eta_{\beta}(l)|\ \|(\mathcal{N}+5)^{n+1}\psi\|.$ (5.38)
@@
$\|I_{3}\psi\|$ $\leq C\|\eta_{s}\|\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}}\int_{0}^{1}\textnormal{d}\tau\ \|\left(\|\eta_{s}\|(\mathcal{N}+1)^{1/2}+C\|\eta_{s}\|^{2}\right)W_{1-\tau}(\eta_{s})\mathcal{R}^{1-\tau}_{\alpha}(k)W_{\tau}(\eta_{s})\psi\|$
$\leq C\|\eta_{s}\|^{2}\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}}\int_{0}^{1}\textnormal{d}\tau\ \|\mathcal{N}W_{1-\tau}(\eta_{s})\mathcal{R}^{1-\tau}_{\alpha}(k)W_{\tau}(\eta_{s})\psi\|+C\|\eta_{s}\|^{2}\|I_{2}\psi\|$
$\leq C\|\eta_{s}\|^{2}\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}}\int_{0}^{1}\textnormal{d}\tau\ e^{C\|\eta_{s}\|(1-\tau)}\|\mathcal{N}\mathcal{R}^{1-\tau}_{\alpha}(k)W_{\tau}(\eta_{s})\psi\|+C\|\eta_{s}\|^{2}\|I_{2}\psi\|$
@@
$=\int_{\tau}^{1}\textnormal{d}\sigma\ \partial_{y_{i}}e^{(1-\sigma)B}\left(\sum_{l\in\Gamma^{\textnormal{nor}}}\eta_{\alpha}(l)\mathcal{E}_{\alpha}(l,k)\right)e^{\sigma B}+\int_{\tau}^{1}\textnormal{d}\sigma\ e^{(1-\sigma)B}\left(\sum_{l\in\Gamma^{\textnormal{nor}}}ik_{i}\eta_{\alpha}(l)\mathcal{E}_{\alpha}(l,k)\right)e^{\sigma B}$
$\hphantom{=}+\int_{\tau}^{1}\textnormal{d}\sigma\ e^{(1-\sigma)B}\left(\sum_{l\in\Gamma^{\textnormal{nor}}}\eta_{\alpha}(l)\mathcal{E}_{\alpha}(l,k)\right)\partial_{y_{i}}e^{\sigma B}$ (5.40)
@@
×$\times e^{(1-\sigma)B}\left(\sum_{l\in\Gamma^{\textnormal{nor}}}\eta_{\alpha}(l)\mathcal{E}_{\alpha}(l,k)\right)e^{\sigma B},$ (5.41)
$I_{4,2}$ $=2i\int_{\tau}^{1}\textnormal{d}\sigma\int_{0}^{1-\sigma}\textnormal{d}a\ e^{(1-a)B}\text{Im}\langle ik_{i}\eta_{s},\mathcal{R}^{1-\sigma-a}\rangle_{k,\alpha}e^{aB}\left(\sum_{l\in\Gamma^{\textnormal{nor}}}\eta_{\alpha}(l)\mathcal{E}_{\alpha}(l,k)\right)e^{\sigma B},$ (5.42)
$I_{4,3}$ $=\int_{\tau}^{1}\textnormal{d}\sigma\ e^{(1-\sigma)B}\left(\sum_{l\in\Gamma^{\textnormal{nor}}}ik_{i}\eta_{\alpha}(l)\mathcal{E}_{\alpha}(l,k)\right)e^{\sigma B},$ (5.43)
$I_{4,4}$ $=\int_{\tau}^{1}\textnormal{d}\sigma\ e^{(1-\sigma)B}\left(\sum_{l\in\Gamma^{\textnormal{nor}}}\eta_{\alpha}(l)\mathcal{E}_{\alpha}(l,k)\right)\sigma\left(c^{\ast}(ik_{i}\eta_{s})-c(ik_{i}\eta_{s})+i\sigma\text{Im}\langle\eta_{s},ik_{i}\eta_{s}\rangle\right)e^{\sigma B},$ (5.44)
$I_{4,5}$ $=2i\int_{\tau}^{1}\textnormal{d}\sigma\ e^{(1-\sigma)B}\left(\sum_{l\in\Gamma^{\textnormal{nor}}}\eta_{\alpha}(l)\mathcal{E}_{\alpha}(l,k)\right)\int_{0}^{\sigma}\textnormal{d}a\ e^{(1-a)B}\text{Im}\langle ik_{i}\eta_{s},\mathcal{R}^{\sigma-a}\rangle_{k,\alpha}e^{aB}.$ (5.45)
@@
$\leq C\sum_{l\in\Gamma^{\textnormal{nor}}}|\eta_{\alpha}(l)|\int_{\tau}^{1}\textnormal{d}\sigma\ (1-\sigma)\|\left(\|\eta_{s}\|\ \|(\mathcal{N}+1)e^{(1-\sigma)B}\mathcal{E}_{\alpha}(l,k)e^{\sigma B}\psi\|+C\|\eta_{s}\|^{2}\|\mathcal{E}_{\alpha}(l,k)e^{\sigma B}\psi\|\right)$
$\leq C\|\eta_{s}\|\sum_{l\in\Gamma^{\textnormal{nor}}}|\eta_{\alpha}(l)|\int_{\tau}^{1}\textnormal{d}\sigma\ (1-\sigma)\|\left(e^{C\|\eta_{s}\|(1-\sigma)}\|(\mathcal{N}+3)\mathcal{E}_{\alpha}(l,k)e^{\sigma B}\psi\|+C\|\eta_{s}\|\,\|\mathcal{E}_{\alpha}(l,k)e^{\sigma B}\psi\|\right)$
$\leq C\|\eta_{s}\|MN^{-\frac{2}{3}+\delta}\sum_{l\in\Gamma^{\textnormal{nor}}}|\eta_{\alpha}(l)|\int_{\tau}^{1}\textnormal{d}\sigma\ \left((1-\sigma)e^{C\|\eta_{s}\|(1-\sigma)}+C\|\eta_{s}\|\right)\|(\mathcal{N}+3)^{2}e^{\sigma B}\psi\|$
$\leq C\|\eta_{s}\|MN^{-\frac{2}{3}+\delta}\sum_{l\in\Gamma^{\textnormal{nor}}}|\eta_{\alpha}(l)|\int_{\tau}^{1}\textnormal{d}\sigma\ \left((1-\sigma)e^{C\|\eta_{s}\|(1-\sigma)}+C\|\eta_{s}\|\right)e^{\tilde{C}\|\eta_{s}\|\sigma}\|(\mathcal{N}+5)^{2}\psi\|$ (5.46)
$\leq CMN^{-\frac{2}{3}+\delta}\sum_{l\in\Gamma^{\textnormal{nor}}}|\eta_{\alpha}(l)|\left((1-\tau)e^{C\|\eta_{s}\|(1-\tau)}+C\|\eta_{s}\|e^{C\|\eta_{s}\|}\right)\|(\mathcal{N}+5)^{2}\psi\|$ (5.47)
where we used Lemma A.4 and Lemma 5.1 in the first inequality, Lemma A.3 in the third inequality and Proposition 2.6 in the second and forth inequality. Therefore using $\int_{0}^{1}(1-\tau)e^{y(1-\tau)}\textnormal{d}\tau=y^{-2}\left((y-1)e^{y}+1\right)$ yields
$\int_{0}^{1}\textnormal{d}\tau\ \|\eta_{s}\|\sqrt{\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}\cap\mathcal{I}_{l}}\|I_{4,1}\psi\|^{2}}$ (5.48)
@@
where we used $\sum_{\alpha\in\mathcal{I}_{l}}\bigg{(}\sum_{l\in\Gamma^{\textnormal{nor}}}|\eta_{\alpha}(l)|\bigg{)}^{2}\leq\sum_{l^{\prime},l\in\Gamma^{\textnormal{nor}}}\|\eta(l)\|_{l^{2}}\|\eta(l^{\prime})\|_{l^{2}}\leq C\|\eta\|^{2}$ by Lemma 5.1.

### v2

@@ 3.2 Effective coherent state
$\leq 2\int_{0}^{t}\textnormal{d}s\int_{0}^{1}\textnormal{d}\tau\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}\bigg{(}\|\overline{(h_{y})_{\alpha}(k)}\mathcal{R}^{1-\tau}_{\alpha}(k)e^{\tau B}\psi\|+\|\mathcal{R}^{1-\tau}_{\alpha}(k)^{\ast}(h_{y})_{\alpha}(k)e^{\tau B}\psi\|\bigg{)}$
$\leq 2\lambda\int_{0}^{t}\textnormal{d}s\int_{0}^{1}\textnormal{d}\tau\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}|\hat{V}(k)n_{\alpha}(k)|\bigg{(}\|\mathcal{R}^{1-\tau}_{\alpha}(k)e^{\tau B}\psi\|+\|\mathcal{R}^{1-\tau}_{\alpha}(k)^{\ast}e^{\tau B}\psi\|\bigg{)}$
$\leq C\int_{0}^{t}\textnormal{d}s\int_{0}^{1}\textnormal{d}\tau\ \lambda k_{\textnormal{F}}\|\hat{V}\|_{2}\bigg{(}\bigg{(}\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}\|\mathcal{R}^{1-\tau}_{\alpha}(k)e^{\tau B}\psi\|^{2}\bigg{)}^{1/2}+\bigg{(}\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}\|\mathcal{R}^{1-\tau}_{\alpha}(k)^{\ast}e^{\tau B}\psi\|^{2}\bigg{)}^{1/2}\bigg{)}$
@@
$\leq C(f_{C}(\lambda k_{\textnormal{F}}t)-1)\lambda k_{\textnormal{F}}tMN^{-\frac{2}{3}+\delta}\|(\mathcal{N}+5)\psi\|.$ (5.36)
Again, by using $\psi=\phi\otimes\Omega$ we obtain the desired bound.
With the subsequent Lemma 5.9, we can conclude with a bound on $h_{0}=-\beta\Delta_{y}$ of the form
@@
$\leq C\beta t\left\{\log\left[f_{1}(\lambda k_{\textnormal{F}}t)\right]+\log\left[f_{1}(\lambda k_{\textnormal{F}}t)\right]^{4}\right\}\left\{f_{C}(\lambda k_{\textnormal{F}}t)+1\right\}$ (5.37)
@@
In total we expect, that all terms can be bounded by assumption on the initial condition on $\phi$. We will first focus on the term $\Delta_{y}W(\eta_{s})$. Recall that it holds
$W(\eta_{s})^{\ast}\partial_{y_{i}}W(\eta_{s})=\int_{0}^{1}\textnormal{d}\tau\ e^{-\tau B}\left(\partial_{y_{i}}B_{s}\right)e^{\tau B}$ (5.39)
due to the same calculation as in the proof of Lemma 5.5. With
@@
$\hphantom{=}+2i\int_{0}^{1}\textnormal{d}\tau\ W_{1-\tau}(\eta_{s})\text{Im}\langle ik_{i}\eta_{s},\mathcal{R}^{1-\tau}\rangle_{\Gamma}W_{\tau}(\eta_{s}).$ (5.42)
@@
$\hphantom{=}+2i\int_{0}^{\tau}\textnormal{d}\sigma\ W_{1-\sigma}(\eta_{s})\text{Im}\langle ik_{i}\eta_{s},\mathcal{R}^{\tau-\sigma}\rangle_{\Gamma}W_{\sigma}(\eta_{s})$ (5.43)
@@
$=$ $\left(c^{\ast}(k^{2}\eta_{s})-c(k^{2}\eta_{s})+i\text{Im}\langle\eta_{s},k^{2}\eta_{s}\rangle\right)W(\eta_{s})+2i\int_{0}^{1}\textnormal{d}\tau\ W_{1-\tau}(\eta_{s})\text{Im}\langle k^{2}\eta_{s},\mathcal{R}^{1-\tau}\rangle_{\Gamma}W_{\tau}(\eta_{s})$
@@
$+2i\sum_{i=1}^{3}\int_{0}^{1}\textnormal{d}\tau\ \text{Im}\langle ik_{i}\eta_{s},\partial_{y_{i}}\left\{W_{1-\tau}(\eta_{s})\mathcal{R}^{1-\tau}_{\alpha}(k)W_{\tau}(\eta_{s})\right\}\rangle_{\Gamma}$ (5.44)
$=$ $\left(c^{\ast}(k^{2}\eta_{s})-c(k^{2}\eta_{s})+i\text{Im}\langle\eta_{s},k^{2}\eta_{s}\rangle\right)W(\eta_{s})+2i\int_{0}^{1}\textnormal{d}\tau\ \text{Im}\langle k^{2}\eta_{s},\mathcal{R}^{1-\tau}\rangle_{\Gamma}W_{\tau}(\eta_{s})$
@@
×$\times\left\{\left(c^{\ast}(ik_{i}\eta_{s})-c(ik_{i}\eta_{s})+i\text{Im}\langle\eta_{s},ik_{i}\eta_{s}\rangle\right)W(\eta_{s})+2i\int_{0}^{1}\textnormal{d}\tau\ \text{Im}\langle ik_{i}\eta_{s},\mathcal{R}^{1-\tau}\rangle_{\Gamma}W_{\tau}(\eta_{s})\right\}$
$+2i\sum_{i=1}^{3}\int_{0}^{1}\textnormal{d}\tau\ \text{Im}\langle ik_{i}\eta_{s},\partial_{y_{i}}\left\{W_{1-\tau}(\eta_{s})\mathcal{R}^{1-\tau}_{\alpha}(k)W_{\tau}(\eta_{s})\right\}\rangle_{\Gamma}$ (5.45)
@@
$I_{2}$ $\coloneqq 2i\int_{0}^{1}\textnormal{d}\tau\ W_{1-\tau}(\eta_{s})\text{Im}\langle k^{2}\eta_{s},\mathcal{R}^{1-\tau}\rangle_{\Gamma}W_{\tau}(\eta_{s}),$ (5.47)
$I_{3}$ $\coloneqq 2\sum_{i=1}^{3}\left(c^{\ast}(ik_{i}\eta_{s})-c(ik_{i}\eta_{s})+i\text{Im}\langle\eta_{s},ik_{i}\eta_{s}\rangle\right)\int_{0}^{1}\textnormal{d}\tau\ iW_{1-\tau}(\eta_{s})\text{Im}\langle ik_{i}\eta_{s},\mathcal{R}^{1-\tau}\rangle_{\Gamma}W_{\tau}(\eta_{s}),$ (5.48)
$I_{4}$ $\coloneqq 2i\sum_{i=1}^{3}\int_{0}^{1}\textnormal{d}\tau\ \text{Im}\langle ik_{i}\eta_{s},\partial_{y_{i}}\left\{W_{1-\tau}(\eta_{s})\mathcal{R}^{1-\tau}_{\alpha}(k)W_{\tau}(\eta_{s})\right\}\rangle_{\Gamma}.$ (5.49)
@@
For $I_{1}$, we can treat all $c^{\ast}(\cdots)$ and $c(\cdots)$ terms with Lemma 5.6 and Lemma 5.8. Furthermore we use that $c_{\alpha}(k)\mathcal{N}=(\mathcal{N}+2)c_{\alpha}(k)$ to estimate
@@
For $I_{3}$, we first observe that for $n\in\mathbb{N}$ and $\mu\in[0,1]$
$\|\mathcal{N}^{n}\mathcal{R}^{\mu}_{\beta}(k)\psi\|$ $\leq\int_{0}^{\mu}\textnormal{d}\tau\ \|\mathcal{N}^{n}e^{-\tau B}\left(\sum_{l\in\Gamma}\eta_{\beta}(l)\mathcal{E}_{\beta}(l,k)\right)e^{\tau B}\psi\|$
$\leq\int_{0}^{\mu}\textnormal{d}\tau\ e^{C\|\eta_{s}\|\tau}\sum_{l\in\Gamma}|\eta_{\beta}(l)|\ \|\mathcal{E}_{\beta}(l,k)(\mathcal{N}+3)^{n}e^{\tau B}\psi\|$
$\leq C\int_{0}^{\mu}\textnormal{d}\tau\ e^{C\|\eta_{s}\|\tau}\sum_{l\in\Gamma}|\eta_{\beta}(l)|\ MN^{-\frac{2}{3}+\delta}\|\mathcal{N}(\mathcal{N}+3)^{n}e^{\tau B}\psi\|$
$\leq C\|\eta_{s}\|^{-1}(e^{C\|\eta_{s}\|\mu}-1)MN^{-\frac{2}{3}+\delta}\sum_{l\in\Gamma}|\eta_{\beta}(l)|\ \|(\mathcal{N}+5)^{n+1}\psi\|.$ (5.52)
@@
$\|I_{3}\psi\|$ $\leq C\|\eta_{s}\|\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}\int_{0}^{1}\textnormal{d}\tau\ \|\left(\|\eta_{s}\|(\mathcal{N}+1)^{1/2}+C\|\eta_{s}\|^{2}\right)W_{1-\tau}(\eta_{s})\mathcal{R}^{1-\tau}_{\alpha}(k)W_{\tau}(\eta_{s})\psi\|$
$\leq C\|\eta_{s}\|^{2}\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}\int_{0}^{1}\textnormal{d}\tau\ \|\mathcal{N}W_{1-\tau}(\eta_{s})\mathcal{R}^{1-\tau}_{\alpha}(k)W_{\tau}(\eta_{s})\psi\|+C\|\eta_{s}\|^{2}\|I_{2}\psi\|$
$\leq C\|\eta_{s}\|^{2}\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}\int_{0}^{1}\textnormal{d}\tau\ e^{C\|\eta_{s}\|(1-\tau)}\|\mathcal{N}\mathcal{R}^{1-\tau}_{\alpha}(k)W_{\tau}(\eta_{s})\psi\|+C\|\eta_{s}\|^{2}\|I_{2}\psi\|$
@@
$=\int_{\tau}^{1}\textnormal{d}\sigma\ \partial_{y_{i}}e^{(1-\sigma)B}\left(\sum_{l\in\Gamma}\eta_{\alpha}(l)\mathcal{E}_{\alpha}(l,k)\right)e^{\sigma B}+\int_{\tau}^{1}\textnormal{d}\sigma\ e^{(1-\sigma)B}\left(\sum_{l\in\Gamma}ik_{i}\eta_{\alpha}(l)\mathcal{E}_{\alpha}(l,k)\right)e^{\sigma B}$
$\hphantom{=}+\int_{\tau}^{1}\textnormal{d}\sigma\ e^{(1-\sigma)B}\left(\sum_{l\in\Gamma}\eta_{\alpha}(l)\mathcal{E}_{\alpha}(l,k)\right)\partial_{y_{i}}e^{\sigma B}$ (5.54)
@@
[…]
$I_{4,2}$ $=2i\int_{\tau}^{1}\textnormal{d}\sigma\int_{0}^{1-\sigma}\textnormal{d}a\ e^{(1-a)B}\text{Im}\langle ik_{i}\eta_{s},\mathcal{R}^{1-\sigma-a}\rangle_{\Gamma}e^{aB}\left(\sum_{l\in\Gamma}\eta_{\alpha}(l)\mathcal{E}_{\alpha}(l,k)\right)e^{\sigma B},$ (5.56)
$I_{4,3}$ $=\int_{\tau}^{1}\textnormal{d}\sigma\ e^{(1-\sigma)B}\left(\sum_{l\in\Gamma}ik_{i}\eta_{\alpha}(l)\mathcal{E}_{\alpha}(l,k)\right)e^{\sigma B},$ (5.57)
$I_{4,4}$ $=\int_{\tau}^{1}\textnormal{d}\sigma\ e^{(1-\sigma)B}\left(\sum_{l\in\Gamma}\eta_{\alpha}(l)\mathcal{E}_{\alpha}(l,k)\right)\sigma\left(c^{\ast}(ik_{i}\eta_{s})-c(ik_{i}\eta_{s})+i\sigma\text{Im}\langle\eta_{s},ik_{i}\eta_{s}\rangle\right)e^{\sigma B},$ (5.58)
$I_{4,5}$ $=2i\int_{\tau}^{1}\textnormal{d}\sigma\ e^{(1-\sigma)B}\left(\sum_{l\in\Gamma}\eta_{\alpha}(l)\mathcal{E}_{\alpha}(l,k)\right)\int_{0}^{\sigma}\textnormal{d}a\ e^{(1-a)B}\text{Im}\langle ik_{i}\eta_{s},\mathcal{R}^{\sigma-a}\rangle_{\Gamma}e^{aB}.$ (5.59)
@@
$\leq C\sum_{l\in\Gamma}|\eta_{\alpha}(l)|\int_{\tau}^{1}\textnormal{d}\sigma\ (1-\sigma)\left(\|\eta_{s}\|\ \|(\mathcal{N}+1)e^{(1-\sigma)B}\mathcal{E}_{\alpha}(l,k)e^{\sigma B}\psi\|+C\|\eta_{s}\|^{2}\|\mathcal{E}_{\alpha}(l,k)e^{\sigma B}\psi\|\right)$
$\leq C\|\eta_{s}\|\sum_{l\in\Gamma}|\eta_{\alpha}(l)|\int_{\tau}^{1}\textnormal{d}\sigma\ (1-\sigma)\left(e^{C\|\eta_{s}\|(1-\sigma)}\|(\mathcal{N}+3)\mathcal{E}_{\alpha}(l,k)e^{\sigma B}\psi\|+C\|\eta_{s}\|\,\|\mathcal{E}_{\alpha}(l,k)e^{\sigma B}\psi\|\right)$
$\leq C\|\eta_{s}\|MN^{-\frac{2}{3}+\delta}\sum_{l\in\Gamma}|\eta_{\alpha}(l)|\int_{\tau}^{1}\textnormal{d}\sigma\ \left((1-\sigma)e^{C\|\eta_{s}\|(1-\sigma)}+C\|\eta_{s}\|\right)\|(\mathcal{N}+3)^{2}e^{\sigma B}\psi\|$
$\leq C\|\eta_{s}\|MN^{-\frac{2}{3}+\delta}\sum_{l\in\Gamma}|\eta_{\alpha}(l)|\int_{\tau}^{1}\textnormal{d}\sigma\ \left((1-\sigma)e^{C\|\eta_{s}\|(1-\sigma)}+C\|\eta_{s}\|\right)e^{\tilde{C}\|\eta_{s}\|\sigma}\|(\mathcal{N}+5)^{2}\psi\|$ (5.60)
$\leq CMN^{-\frac{2}{3}+\delta}\sum_{l\in\Gamma}|\eta_{\alpha}(l)|\left((1-\tau)e^{C\|\eta_{s}\|(1-\tau)}+C\|\eta_{s}\|e^{C\|\eta_{s}\|}\right)\|(\mathcal{N}+5)^{2}\psi\|$ (5.61)
where we used Lemma A.4, Lemma 5.6 and Lemma 5.8 in the first inequality, Lemma A.3 in the third inequality and Proposition 5.4 in the second and forth inequality. Therefore using $\int_{0}^{1}(1-\tau)e^{y(1-\tau)}\textnormal{d}\tau=y^{-2}\left((y-1)e^{y}+1\right)$ yields
$\int_{0}^{1}\textnormal{d}\tau\ \|\eta_{s}\|\sqrt{\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}\cap\mathcal{I}_{l}}\|I_{4,1}\psi\|^{2}}$ (5.62)
@@
where we used $\sum_{\alpha\in\mathcal{I}_{l}}\bigg{(}\sum_{l\in\Gamma}|\eta_{\alpha}(l)|\bigg{)}^{2}\leq\sum_{l^{\prime},l\in\Gamma}\|\eta(l)\|_{l^{2}}\|\eta(l^{\prime})\|_{l^{2}}\leq C\|\eta\|^{2}$ by Lemma 5.6 and Lemma 5.8.